OpenStudy (anonymous):

The equation, in general form, of the line that passes through the point (3,8) and is parallel to the line 5x−8y+9=0 is Ax+By+C=0

OpenStudy (anonymous):

my answer isnt matching up...i got -5x+8y-49=0

OpenStudy (anonymous):

I have found 5x-8y+49=O

OpenStudy (anonymous):

if they are parallel that means the have the same slope, so you can solve for the slope and then apply it to solve for the b value of your other line :)

OpenStudy (anonymous):

i did...but how did you get your signs different than mine _jacobian

OpenStudy (anonymous):

Write the given equation in the standard format\[y=mx+b where m is the slope and b is y intercept.\]

OpenStudy (anonymous):

5x-8y+9 = y=5/8+9/8

OpenStudy (anonymous):

slope=5/8 plug in your point 3,8 8=5/8(3)+b b=49/8 y=5/8x+49/8

OpenStudy (anonymous):

get it :)?

OpenStudy (anonymous):

I agree Mythias... :)

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

sry for the slowness, my computer is majorly lagging LOL

OpenStudy (anonymous):

\[y=5/8 x + 49/8\] slope m = 5/8. use and (x1, y1)=(3,8) in \[y-y1=m(x-x1)\]